# Language Models That Walk the Talk: A Framework for Formal Fairness Certificates

### v1

@@ 4.2 Main results and discussions
In this section, we present the results of our experiments. We evaluated our approach on text classification language models of up to 10 blocks of transformers. We provide details on the pre-training hyperparameters in Section C.1 and ablation studies including more models in Section C.3.
Figure 2: Verifiability radar plot with gender-distance reference circle for 8-block transformer in gender fairness mitigation.
Figure 3: Verifiability radar plot with gender-distance reference circle for 10-block transformer in gender fairness mitigation.
Table 1: Model performance and verification metrics.
Model Salary Prediction & Gender Bias Mitigation
Test Acc. ($\uparrow$) Fairness Score $\Psi$ ($\uparrow$)
8 blocks 88% 100%
10 blocks 88% 20%
Model Toxicity Detection & Content Moderation
Test Acc. ($\uparrow$) Fairness Score $\Psi$ ($\uparrow$)
3 blocks 91% 75%
@@
We evaluate the fairness of our models using the fairness score $\Psi$ defined as the percentage of sentences in which all gender-related tokens fall within a semantic safety threshold (Section 3.3). These tokens are identified as the 50 nearest neighbors to the words male and female in the embedding space, as these are the primary gender indicators in the Adult Census dataset. Notably, male and female appear within each other’s nearest neighbors and both share the maximum $\ell_{\infty}$ distance of $3\times 10^{-4}$, which defines the reference radius (shown as the coral circle in Figures 3 and 3). Tokens within this radius are considered semantically equal (Table 2). A complete list of gender-related neighbors is provided in Section C.3.4.
The radar plots in Figures 3 and 3 illustrate per-sentence verifiability for the salary prediction language model configured with 8 and 10 transformer blocks, respectively. Each axis corresponds to a sentence, and the tip of the blue region represents its maximum verifiable perturbation radius $\epsilon_{\text{max}}$. Sentences with blue tips outside the coral-colored dashed reference circle (marking the largest $\ell_{\infty}$ distance $\delta$) are considered verified (i.e., fair), while red dots highlight those sentences that could not be verified. Table 1 presents the model’s test accuracy alongside the corresponding fairness scores after training.
@@
It is important to emphasize that the primary objective of this work is not solely to ensure that every evaluated model is verified as fair or safe, but rather to establish a reliable framework for assessing these properties. Ideally, our framework exposes adversarial examples and biases and can guide the development of fairer systems, such as through pre-training strategies that cluster gendered tokens into the same semantic class. Thus, when the verification results reveal unfairness, such as in the case where the 10-block model was found to be less fair than its 8-block counterpart, this indicates that the model might need further training towards fairer predictions.
Figure 4: Verifiability radar plot with toxicity-distance reference circle for the 3-block toxicity detection model in content moderation.
@@
Toxicity detection.
For the toxicity detection experiment, we begin by identifying a set of common toxicity-related terms from the dataset. Using a pre-trained embedding space, we then retrieve the 50 nearest neighbors for each term, following the same approach used for gendered terms. For each toxic word, we calculate the maximum $\ell_{\infty}$ distance among its top 50 nearest neighbors. This value represents the farthest semantic deviation closely related within the embedding space to still be considered as having the same semantic meaning. The largest such distance observed across all common toxic terms defines the reference circle used in this experiment (shown as the coral circle in Figure 4). An example of the full neighbor set is provided in Section C.3.4.
The per-sentence verifiability results for the 3-block transformer language model used in toxicity detection are presented in Figure 4. Table 1 presents the model’s test accuracy alongside the corresponding fairness scores after training. In addition, Figure 4 presents the pairwise $\ell_{\infty}$ distances between frequently occurring toxicity-related word pairs in the Jigsaw dataset and throughout verification [^3] It is worth noting that the distances between toxic words after pre-training tend to be larger than those observed among gendered word pairs. This can be attributed to the noisier nature of the toxicity-related dataset, which introduces greater variability and makes it more challenging to learn embedding distances.. The majority of them fall within the maximum verifiable noise radius for every sentence. This is particularly important in the context of content moderation. For example, an Internet user might attempt to evade detection by slightly modifying toxic terms (e.g., changing arse to azzhole). However, our model would still be able to verify such edits and maintain detection capability, demonstrating the effectiveness of our verification approach in real-world scenarios such as toxicity detection and content moderation [^4] The authors opted to use a relatively small transformer model for the toxicity detection task. Unlike salary prediction, which requires learning complex relationships (e.g., between demographic attributes like gender, race, native countries, and income level), toxicity detection is a comparatively simpler classification problem. The presence of explicit toxicity-related words makes it easier for the model to identify toxic content. Therefore, increasing the number of transformer blocks was deemed unnecessary and would likely lead to overfitting—a concern that was also observed empirically..
@@ 5 Related work
This work continues the progressive effort to formally verify neural networks in recent years (brix2023fourth, ): As neural networks are vulnerable to adversarial attacks (goodfellow2015explaining, ), formally verifying them against input perturbations is of immense importance in safety-critical scenarios. Neural network verifiers can generally be categorized into optimization-based verifiers zhang2018efficient ; katz2019marabou ; henriksen2020efficient ; singh2019abstract and approaches using set-based computing gehr2018ai2 ; singh2018fast ; lopez2023nnv ; kochdumper2022open . Early approaches focused on computing the exact output set of neural networks with ReLU activations, which has been shown to be NP-hard katz2017reluplex . Convex relaxations of the problem singh2018fast ; xu2020fast and efficient branch-and-bound techniques wang2021beta have shown to achieve impressive verification results brix2023fourth . Recently, non-convex relaxations have also been shown to be effective in this field of research kochdumper2022open ; wang2023polar ; ladner2023automatic ; fatnassi2023bern ; ortiz2023hybrid .
@@
Many works have focused on verifying the safety and fairness of neural networks. For example, FairNNV tumlin2024fairnnv verifies individual and counterfactual fairness via reachability analysis, while safety in robotic systems has been studied using formal methods sun2019formal . Probabilistic approaches, such as those using Markov Chains sun2021probabilistic , FairSquare 10.1145/3133904 , and VeriFair 10.1145/3360544 , enable rigorous group fairness verification. The OVERT algorithm sidrane2022overt enables sound safety verification of nonlinear, discrete-time, closed-loop dynamical systems with neural network control policies. Discriminatory input detection and fairness analysis are addressed by DeepGemini xie2023deepgemini , complemented by Fairify’s SMT-based method for verifying individual fairness biswas2023fairify . An automated verification framework for feed-forward multi-layer neural networks based on Satisfiability Modulo Theories (SMT) is proposed in huang2017safety , while NNV tran2020nnv offers a set-based verification approach tailored for deep neural networks. Other methods include training-based fairness improvement borcatasciuc2022provablefairnessneuralnetwork , certified representation learning ruoss2020learning , structured data verification john2020verifying , global fairness verification with CertiFair khedr2023certifair , parallel static analysis for causal fairness urban2020perfectly , and Bayesian and satisfiability-based tools like FVGM and Justicia ghosh2021justicia ; ghosh2022algorithmic . A new approach to measuring and addressing discrimination in supervised learning, where a model is trained to predict an outcome based on input features hardt2016equality . Prior work has predominantly focused on fairness verification of standard feed-forward neural networks, whereas our approach focuses on verifying both the fairness and safety properties of transformer-based language models.
@@
Most bias evaluation metrics for large language models can be categorized by what they use from the model gallegos2024bias , such as the embeddings caliskan2017semantics ; dev2020measuring ; guo2021detecting , model-assigned probabilities webster2020measuring ; kurita-etal-2019-measuring ; salazar-etal-2020-masked ; wang-cho-2019-bert ; nadeem-etal-2021-stereoset ; kaneko2022unmasking or generated text rajpurkar-etal-2016-squad ; bordia-bowman-2019-identifying ; liang2023holistic ; gehman-etal-2020-realtoxicityprompts ; sicilia-alikhani-2023-learning ; nozza-etal-2021-honest ; 10.1145/3442188.3445924 . Here, we focus on the generated text, given that it is the primary focus of our study. Language models’ responses to prompts can reveal social biases cheng-etal-2023-marked , using tasks like prompt completion, conversation, and analogical reasoning li2024surveyfairnesslargelanguage . The standard method involves prompting the model and evaluating its continuation for bias gallegos2024bias . GPT-4 has been shown to generate more racially stereotyped personas than human writers when prompted to describe demographic groups cheng-etal-2023-marked . Similarly, earlier work validated gender bias in GPT-3’s output using 800 prompts (e.g., “[He] was very” vs. “[She] was very”), revealing stereotypes in the model output and highlighting the need for debiasing brown2020language ; 10.1145/3461702.3462624 . Other analyses show that language models frequently associate certain occupations with specific genders based on co-occurrence patterns liang2023holistic . The fairness of a large language model is also sometimes evaluated using a second, more powerful, large language model eloundou2024first , which does not provide formal guarantees. Rather, instead of finding the vulnerability of the first model, it can try to fool the second model to “pass” the fairness test wang-etal-2024-large-language-models-fair .
@@
While the vulnerability of large language models to adversarial prompts has been explored sun2024trustllm , research on the non-existence of such prompts remains limited, and the formal verification of transformer-based language models remains largely unexplored. As with standard neural networks, the method proposed by shi2020robustness verifies robustness using linear bounds, allowing for at most two perturbed words. This work has been extended to the zonotope domain (bonaert2021fast, ), enabling the verification of longer sentences with all words being perturbed. Other works also investigated finding better bounds of the softmax function within transformers (wei2023convex, ; shi2024neural, ) and investigated the preservation of dependencies between attention heads (zhang2024galileo, ).
@@
Existing efforts—including this work—focus on perturbations applied directly to the embedding space of the input in large language models vaswani2017attention . Please note that most existing works apply arbitrary $\epsilon$-perturbations to the input and examine whether such changes can alter the predictions of language models. However, they often overlook a critical assumption: Whether semantically meaningful changes, such as synonym substitution or variation across gender or toxicity-related terms, are actually captured by the $\epsilon$-ball. When synonyms are considered, the focus tends to be limited to traditional definitions, without exploring how the embedding space could be shaped to support other important criteria, such as mitigating gender bias or detecting adversarially crafted yet semantically equivalent toxic terms. In this work, we address this gap by offering a holistic perspective on the formal verification of language models.
@@ 6 Limitation
Given the scale and complexity of modern large language models achiam2024gpt4 , existing verification techniques face a fundamental trade-off between precision and computational cost, rendering them impractical for large-scale models shi2020robustness ; bonaert2021fast . Consequently, we exclude larger pre-trained models such as BERT from our evaluation due to their verification intractability.
Furthermore, our study uses a single state-of-the-art verification method based on zonotopes, applied exclusively to binary text classification models. This method is sound but incomplete; thus, while an obtained fairness certificate always holds, the method may fail to verify models that are, in fact, fair. Nonetheless, our verification framework is agnostic to the underlying verification method, and we hope that with the continuous progress in the field of neural network verification brix2023yearsinternationalverificationneural , The fairness of more complex architectures can be verified using our framework.
@@ 7 Conclusion
We present a generalized zonotope-based verification framework for formally assessing the fairness and safety of language models and demonstrate it on models with up to 10 transformer blocks. In particular, the framework consists of carefully designing the embedding space to capture synonyms, gender-related terms, or toxicity-related terms through perturbation using pre-training-based transfer learning. This embedding layer is then integrated into the custom transformer models for downstream language model training, ensuring the fairness of their predictions. Lastly, the fairness of the models is formally verified using an approach based on zonotopes. Importantly, this formal verification technique circumvents the combinatorial complexity of testing individual word substitutions or synonyms by enabling sentence-level verification.
Using this holistic framework, we made sure that gender-related terms are embedded into sufficient proximity through targeted pre-training, allowing different gender terms to be treated as semantically equivalent during inference. This embedding structure is critical in enabling formal verification of gender fairness, as model outputs remained invariant under gender word substitutions. Additionally, we applied the same approach to toxicity-related terms, showing that semantically similar toxic words clustered closely and were consistently detected, even under adversarial perturbations. This highlights the robustness of the system in content moderation tasks. As a result, the framework achieved a fairness score of 100% for an 8-block transformer language model in the salary prediction task, and 75% for a 3-block toxicity detection model. Notably, such results cannot be obtained if the embedding space is not properly designed. These findings underscore the potential utility of our framework in gender bias mitigation for fairness enforcement and its reliability in supporting content moderation systems.
Acknowledgments
This work was supported by the project FAI (No. 286525601) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).

### v2

@@ 4.2 Main results and discussions
In this section, we present the results of our experiments. We evaluated our approach on text classification language models of up to 10 blocks of transformers. We provide details on the pre-training hyperparameters in Section C.1 and ablation studies, including more models in Section C.3.
(a) 8-block model.
(b) 10-block model.
Figure 2: Verifiability radar plot with gender-distance reference circle (coral dashed circle) in gender bias mitigation with 8 and 10 blocks of transformers. The axes S1 to S50 stand for the 50 sentences from the held-out test dataset.
Table 1: Model performance (test accuracy) and verification metrics. ’NA’ (Not Applicable) indicates that gender terms were omitted; hence, fairness could not be computed.
[…]
Model Test Acc. ($\uparrow$) Fairness Score $\Psi$ ($\uparrow$)
Gender attributes excluded
8 blocks 82% NA
10 blocks 83% NA
Gender attributes included
8 blocks 88% 90%
10 blocks 88% 12%
[…]
Model Test Acc. ($\uparrow$) Fairness Score $\Psi$ ($\uparrow$)
3 blocks 95% 50%
@@
We evaluate the fairness of our models using the fairness score $\Psi$ from eq (4). To compute the fairness score, we constructed the test dataset by randomly selecting 50 sentences—25 from each class (label = 0 or 1 for whether the salary exceeds \$50,000)—from the training set. This represents the model’s overall fairness. Gender-related tokens are identified as the 50 closest neighbors to the word male and female (the primary gender indicators in the Adult Census dataset) in the embedding space. The largest $\ell_{\infty}$ distance between any of their nearest 50 neighbors is $3\times 10^{-4}$ (for both the word male and female. Notably, male and female appear within each other’s 50 nearest neighbors, demonstrating the effectiveness of pre-training). We use this value as a reference threshold (illustrated by the coral circle in Figure 2). Tokens within this threshold are treated as semantically equivalent (see Table 2). A full list of gender-related neighbors is provided in Section C.3. Table 1 presents the model’s test accuracy and fairness scores after training. Test accuracy is a crucial component of our evaluation, as fairness assessments are only meaningful when applied to models that demonstrate reliable predictive performance. Figure 2 shows radar plots of per-sentence verifiability, where each axis represents a sentence and the blue region’s tip indicates the maximum verifiable perturbation radius $\epsilon_{\text{max}}$. Sentences with tips outside the coral dashed circle (radius $\delta$) are deemed verifiable (fair), while red dots mark unverifiable cases.
The 8-block language model achieved a fairness score of 90%, indicating that its predictions remain robust and unbiased with respect to gender across all tested scenarios in the salary prediction task involving demographic attributes. We are noting that the fairness scores of the model with a deeper architecture decreased compared with smaller models (10 vs. 8 transformer blocks), despite exhibiting similar test accuracy. This decline might be due to larger models being harder to verify (Katz et al., 2017) or the model indeed being unfair. To this date, we treat both cases the same way and only count models as fair if they are verifiably so. However, it is important to emphasize that the primary objective of this work is not solely to ensure that every evaluated model is verified as fair or safe, but rather to establish a reliable framework for assessing these properties. Ideally, our framework exposes adversarial examples and biases and can guide the development of fairer systems, such as through pre-training strategies that cluster gendered tokens into the same semantic class on the embedding layers utilized in the model. Thus, when the verification results reveal unfairness or potential unfairness, this indicates that the model might need further training towards fairer predictions.
@@
While removing gender attributes from the input may seem like a straightforward way to mitigate bias, this approach is often insufficient in practice. A prior study (Valentim et al., 2019) shows that removing sensitive attributes like gender does not always lead to a fairer model, as models can still learn bias from correlated features—and may even become more biased by relying on proxy attributes. Similar studies support this, showing that simply removing sensitive attributes is insufficient, as protected features may be encoded or correlated with other variables (Loi et al., 2021; González-Sendino et al., 2024; Goethals et al., 2024). Our experimental result (Table 1) further shows that excluding gender-related information in both the pre-training and language model training phase leads to a drop in model accuracy, indicating that such features can play a legitimate role in prediction, highlighting the necessity of our proposed fairness score to assess and manage potential bias rather than outright exclusion.
Figure 3: Verifiability radar plot with toxicity-distance reference circle (coral dashed circle) for the 3-block toxicity detection model in content moderation. The axes S1 to S50 stand for the 50 sentences from the held-out test dataset.
Toxicity detection.
For the toxicity detection experiment, we begin by identifying a set of common toxicity-related terms from the dataset. Using a pre-trained embedding space, we then retrieve the 50 nearest neighbors for each term, following the same approach used for gendered terms. For each toxic word, we compute the maximum $\ell_{\infty}$ distance among its 50 nearest neighbors, representing the semantic boundary for meaning-preserving substitutions. The largest of these distances defines the reference circle (coral, in Figure 3), used to assess whether the model’s predictions remain stable under semantically equivalent toxic term replacements. An example of the full neighbor set is provided in Section C.3.
@@
The per-sentence verifiability results for the 3-block transformer model used in toxicity detection are shown in Figure 3. The model achieved a test accuracy of 95% and a fairness score of 50% after training and formal verification. This indicates that the model can reliably censor toxic content—regardless of any adversarial paraphrasing using toxic equivalents—in 50% of the tested cases. In addition, Table 3 presents the pairwise $\ell_{\infty}$ distances between frequently occurring toxicity-related word pairs in the Jigsaw dataset and throughout verification. The majority of them fall within the maximum verifiable noise radius for every sentence. This is particularly important in the context of content moderation. For example, an Internet user might attempt to evade detection by slightly modifying toxic terms (e.g., changing arse to azzhole). However, a model that has obtained a fairness certificate would guarantee to be able to verify such edits and maintain detection capability, demonstrating the effectiveness of our verification approach in real-world scenarios such as toxicity detection and content moderation [^3] The authors used a smaller transformer for toxicity detection, as explicit cues make the task simpler and larger models led to overfitting, as observed empirically..
To further substantiate the effectiveness of our framework, we include an ablation study in Section C.3 that compares the fairness scores of normally trained and adversarially trained models. The results of this ablation study provide additional evidence of the framework’s ability to distinguish and reward models that incorporate adversarial training.
@@ 5 Related work
This work continues the progressive effort to formally verify neural networks in recent years (Brix et al., 2023a): As neural networks are vulnerable to adversarial attacks (Goodfellow et al., 2015b), formally verifying them against input perturbations is of immense importance in safety-critical scenarios. Neural network verifiers can generally be categorized into optimization-based verifiers (Zhang et al., 2018; Katz et al., 2019; Henriksen and Lomuscio, 2020; Singh et al., 2019) and approaches using set-based computing (Gehr et al., 2018b; Singh et al., 2018; Lopez et al., 2023; Kochdumper et al., 2023). Early approaches focused on computing the exact output set of neural networks with ReLU activations, which has been shown to be NP-hard (Katz et al., 2017). Convex relaxations of the problem (Singh et al., 2018; Xu et al., 2021) and efficient branch-and-bound techniques (Wang et al., 2021) have shown to achieve impressive verification results (Brix et al., 2023a). Recently, non-convex relaxations have also been shown to be effective in this field of research (Kochdumper et al., 2023; Wang et al., 2023; Ladner and Althoff, 2023; Fatnassi et al., 2023; Ortiz et al., 2023).
@@
Many works have focused on verifying the safety and fairness of neural networks. For example, FairNNV (Tumlin et al., 2024) verifies individual and counterfactual fairness via reachability analysis, while safety in robotic systems has been studied using formal methods (Sun et al., 2019). Probabilistic approaches, such as those using Markov Chains (Sun et al., 2021), FairSquare (Albarghouthi et al., 2017), and VeriFair (Bastani et al., 2019), enable rigorous group fairness verification. The OVERT algorithm (Sidrane et al., 2022) enables sound safety verification of nonlinear, discrete-time, closed-loop dynamical systems with neural network control policies. Discriminatory input detection and fairness analysis are addressed by DeepGemini (Xie et al., 2023), complemented by Fairify’s SMT-based method for verifying individual fairness (Biswas and Rajan, 2023). An automated verification framework for feed-forward multi-layer neural networks based on Satisfiability Modulo Theories (SMT) is proposed in (Huang et al., 2017), while NNV (Tran et al., 2020) offers a set-based verification approach tailored for deep neural networks. Other methods include training-based fairness improvement (Borca-Tasciuc et al., 2022), certified representation learning (Ruoss et al., 2020), structured data verification (John et al., 2020), global fairness verification with CertiFair (Khedr and Shoukry, 2023), parallel static analysis for causal fairness (Urban et al., 2020), and Bayesian and satisfiability-based tools like FVGM and Justicia (Ghosh et al., 2021, 2022). A new approach to measuring and addressing discrimination in supervised learning, where a model is trained to predict an outcome based on input features (Hardt et al., 2016). Prior work has predominantly focused on fairness verification of standard feed-forward neural networks, whereas our approach focuses on verifying both the fairness and safety properties of transformer-based language models.
@@
Bias evaluation metrics for large language models can be categorized by what they use from the model (Gallegos et al., 2024), such as the embeddings (Caliskan et al., 2017a; Dev et al., 2020; Guo and Caliskan, 2021), model-assigned probabilities (Webster et al., 2020; Kurita et al., 2019; Salazar et al., 2020; Wang and Cho, 2019; Nadeem et al., 2021; Kaneko and Bollegala, 2022) or generated text (Rajpurkar et al., 2016; Bordia and Bowman, 2019; Liang et al., 2023; Gehman et al., 2020; Sicilia and Alikhani, 2023; Nozza et al., 2021; Dhamala et al., 2021). Here, we focus on the generated text, as it is the primary focus of our study. Language models’ responses to prompts can reveal social biases (Cheng et al., 2023), using tasks like prompt completion, conversation, and analogical reasoning (Li et al., 2024). The standard method involves prompting the model and evaluating its continuation for bias (Gallegos et al., 2024). GPT-4 has been shown to generate more racially stereotyped personas than human writers when prompted to describe demographic groups (Cheng et al., 2023). Similarly, earlier work validated gender bias in GPT-3’s output using 800 prompts (e.g., “[He] was very” vs. “[She] was very”), revealing stereotypes in the model output and highlighting the need for debiasing (Brown et al., 2020; Abid et al., 2021). Other analyses show that language models frequently associate certain occupations with specific genders based on co-occurrence patterns (Liang et al., 2023). The fairness of a large language model is also sometimes evaluated using a second, more powerful, large language model (Eloundou et al., 2024), which does not provide formal guarantees. Rather, instead of finding the vulnerability of the first model, it can try to fool the second model to “pass” the fairness test (Wang et al., 2024).
@@
While the vulnerability of large language models to adversarial prompts has been explored (Sun et al., 2024), research on the non-existence of such prompts remains limited, and the formal verification of transformer-based language models remains largely unexplored. As with standard neural networks, the method proposed by (Shi et al., 2020) verifies robustness using linear bounds, allowing for at most two perturbed words. This work has been extended to the zonotope domain (Bonaert et al., 2021), enabling the verification of longer sentences with all words being perturbed. Other works also investigated finding better bounds of the softmax function within transformers (Wei et al., 2023; Shi et al., 2024) and investigated the preservation of dependencies between attention heads (Zhang et al., 2024).
@@
Existing efforts—including this work—focus on perturbations applied directly to the embedding space of the input in large language models (Vaswani et al., 2017). Please note that most existing works apply arbitrary $\epsilon$-perturbations to the input and examine whether such changes can alter the predictions of language models. However, they often overlook a critical assumption: Whether semantically meaningful changes, such as synonym substitution or variation across gender or toxicity-related terms, are actually captured by the $\epsilon$-ball. When synonyms are considered, the focus tends to be limited to traditional definitions, without exploring how the embedding space could be shaped to support other important criteria, such as mitigating gender bias or detecting adversarially crafted yet semantically equivalent toxic terms. In this work, we address this gap by offering a holistic perspective on the formal verification of language models.
@@ 6 Limitation
Given the scale and complexity of modern large language models (Achiam et al., 2024), existing verification techniques face a fundamental trade-off between precision and computational cost, rendering them impractical for large-scale models (Shi et al., 2020; Bonaert et al., 2021). Consequently, we exclude larger pre-trained models such as BERT from our evaluation due to their verification intractability.
This work focuses on gender as a demographic attribute in salary prediction, but the proposed fairness evaluation framework is attribute-agnostic and can be extended to other variables, such as race. While our experiments address binary classification, the framework is, in principle, applicable to multi-class tasks and broader fairness assessments.
Furthermore, our study uses a single state-of-the-art verification method based on zonotopes, applied exclusively to binary text classification models. This method is sound but incomplete; thus, while an obtained fairness certificate always holds, the method may fail to verify models that are, in fact, fair. Nonetheless, our verification framework is agnostic to the underlying verification method, and we hope that with the continuous progress in the field of neural network verification (Brix et al., 2023b), the fairness of more complex architectures can be verified using our framework.
@@ 7 Conclusion
We present a generalized verification framework for formally assessing the fairness and safety of language models and demonstrate it on models with up to 10 transformer blocks. In particular, the framework consists of carefully designing the embedding space to capture synonyms, gender-related terms, or toxicity-related terms through perturbation using pre-training-based transfer learning. This embedding layer is then integrated into the custom transformer models for downstream language model training, ensuring the fairness of their predictions. Lastly, the fairness of the models is formally verified using an approach based on zonotopes. Importantly, this formal verification technique circumvents the combinatorial complexity of testing individual word substitutions or synonyms by enabling sentence-level verification.
Using this holistic framework, we made sure that gender-related terms are embedded into sufficient proximity through targeted pre-training, allowing different gender terms to be treated as semantically equivalent during inference. This embedding structure is critical in enabling formal verification of gender fairness, as model outputs remained invariant under gender word substitutions. Additionally, we applied the same approach to toxicity-related terms, showing that semantically similar toxic words clustered closely and were consistently detected, even under adversarial perturbations. This highlights the robustness of the system in content moderation tasks. As a result, the framework achieved a fairness score of 90% for an 8-block transformer language model in the salary prediction task, and 50% for a 3-block toxicity detection model. Notably, such results cannot be obtained if the embedding space is not properly designed. These findings underscore the potential utility of our framework in gender bias mitigation for fairness enforcement and its reliability in supporting content moderation systems.
Acknowledgements
This work was supported by the project FAI (No. 286525601) funded by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG).